# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ 2.7. Convex cocompact groups
([17, Prop. 10.1]) Let $\Omega$ be a properly convex domain of $\mathbb{P}(\mathbb{R}^{d})$. For any $x_{0}\in\Omega$, there exists $\kappa>0$ such that for any $g\in\textup{Aut}(\Omega)$,
$\log\frac{\sigma_{1}(g)}{\sigma_{d}(g)}\geqslant 2d_{\Omega}(gx_{0},x_{0})-\kappa.$
Let $\Gamma$ be a subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ preserving a properly convex domain $\Omega$. By using the previous proposition we can control the Gromov product with respect to $\varepsilon_{1}\in\mathfrak{a}^{\ast}$ as follows:
@@
By Theorem 2.5 there exists a finite subset $F$ of $\Gamma$ and $M>0$ such that for every $\gamma\in\Gamma$ there exists $f\in F$ such that $\log\frac{\ell_{1}(\gamma f)}{\ell_{d}(\gamma f)}\geq\log\frac{\sigma_{1}(\gamma))}{\sigma_{d}(\gamma)}-M$. The translation length of an isometry $g\in\textup{Aut}(\Omega)$ is exactly $\frac{1}{2}\log\frac{\ell_{1}(g)}{\ell_{d}(g)}$, see [15, Prop. 2.1]. In particular, if $\gamma\in\Gamma$ and $f\in F$ are as previously, we have that
@@
Then, by Proposition 2.7 and (2.7), we obtain a uniform constant $L>0$ such that
$\ \Big{|}\log\frac{\sigma_{1}(\rho(\gamma))}{\sigma_{d}(\rho(\gamma))}-2d_{\Omega}(\gamma x_{0},x_{0})\Big{|}\leqslant L$
@@
The following lemma follows immediately from [17, Thm. 1.4] and [39, Thm. 1.27] and is used to pass from a $P_{1}$-Anosov representation to a convex cocompact action in some projective space.
@@
Let $V_{d}$ be the vector space of $d\times d$-symmetric matrices and $S_{d}:\mathsf{GL}_{d}(\mathbb{R})\rightarrow\mathsf{GL}(V_{d})$ be the representation defined as follows $S_{d}(g)X=gXg^{t}$ for $X\in V_{d}$. For every $P_{1}$-Anosov representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, the representation $S_{d}\circ\rho$ is $P_{1}$-Anosov and $S_{d}(\rho(\Gamma))$ is a strongly convex cocompact subgroup of $\mathsf{GL}(V_{d})$.
Given two representations $\rho_{1}:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ and $\rho_{2}:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, we say that $\rho_{1}$ uniformly dominates $\rho_{2}$ if there is $\delta\in(0,1)$ with the property that for every $\gamma\in\Gamma$,
$(1-\delta)\log\ell_{1}(\rho_{1}(\gamma))\geqslant\log\ell_{1}(\rho_{2}(\gamma)).$
We will also need the following lemma for the proof of Proposition 1.12, which allows us to control the Cartan projection of an Anosov representation $\rho$ in terms of the Cartan projection of a semisimplification $\rho^{ss}$ of $\rho$. We expect that this fact follows by the techniques of Guichard–Wienhard in [23, §5] showing that Anosov representations have strong proximality properties.
@@
Let $\Gamma$ be a word hyperbolic group, $G$ a real semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$. Suppose $\psi:\Gamma\rightarrow G$ is a $P_{\theta}$-Anosov representation with semisimplification $\psi^{ss}:\Gamma\rightarrow G$. There is a constant $C_{\psi}>0$, depending only on $\psi$, such that for every $\gamma\in\Gamma$,
$\max_{\alpha\in\theta}\Big{|}\big{\langle}\omega_{\alpha},\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big{\rangle}\Big{|}\leqslant C_{\psi}.$
@@
Let us first observe that for any linear representation $\phi:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ and any semisimplification $\phi^{ss}$ of $\phi$, by Theorem 2.5, there exists a constant $Q>0$, depending on $\phi$, such that
$\log\sigma_{1}(\phi(\gamma))\geq\log\sigma_{1}(\phi^{ss}(\gamma))-Q$
for every $\gamma\in\Gamma$. Moreover, for every $\alpha\in\theta$, there exists $N_{\alpha}>0$ such that $N_{\alpha}\omega_{\alpha}$ is the highest weight of an irreducible proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{m}(\mathbb{R})$. In particular, there exists $Q^{\prime}>0$, depending only on $\tau_{\alpha}$, such that $\big{|}\log\sigma_{1}(\tau_{\alpha}(h))-N_{\alpha}\langle\omega_{\alpha},\mu(h)\rangle\big{|}\leq Q^{\prime}$ for every $h\in G$. The previous two facts show that there exists $C_{\rho}>0$, depending only on $\rho$ and $G$, such that for every $\gamma\in\Gamma$,
$\big{\langle}\omega_{\alpha},\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big{\rangle}\geq-C_{\rho}.$
Now we prove that by enlarging $C_{\rho}>0$, for every $\gamma\in\Gamma$ we also have,
$\big{\langle}\omega_{\alpha},\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big{\rangle}\leqslant C_{\rho}.$ (2)
By Proposition 2.4, we may compose $\psi$ with an irreducible representation $\tau:G\rightarrow\mathsf{GL}_{n}(\mathbb{R})$ such that $\rho:=\tau\circ\psi$ and its semisimplification $\rho^{ss}=\tau\circ\psi^{ss}$ are $P_{1}$-Anosov. Up to composing $\rho$ with the representation $S_{n}$ from Lemma 2.10, we may further assume that $\rho(\Gamma)$ and the dual $\rho^{\ast}(\Gamma)$ preserve (possibly different) properly convex domain in $\mathbb{P}(\mathbb{R}^{n})$. Up to conjugating $\rho$, and possibly considering the dual representation of this conjugate, we may assume $\rho(\Gamma)$ preserves a properly convex domain $\Omega_{0}\subset\mathbb{P}(\mathbb{R}^{n})$ and there is a decomposition $\mathbb{R}^{n}=V_{1}\oplus\cdots\oplus V_{\ell}$ such that
$\rho=\begin{pmatrix}\rho_{1}&\ast&\ast\\ &\ddots&\ast\\ &&\rho_{\ell}\end{pmatrix},\ \rho^{ss}=\begin{pmatrix}\rho_{1}&&\\ &\ddots&\\ &&\rho_{\ell}\end{pmatrix}$
where $\rho_{i}:\Gamma\rightarrow\mathsf{GL}(V_{i})$ are irreducible representations, $\rho_{1}$ is $P_{1}$-Anosov and uniformly dominates $\rho_{i}$ for $2\leqslant i\leqslant\ell$. In particular, $\rho_{1}$ is the restriction of $\rho^{ss}$ on the vector subspace $\langle\xi_{\rho^{ss}}\big{(}\partial_{\infty}\Gamma\big{)}\rangle$.
@@
$\rho(\gamma)=\begin{pmatrix}\rho_{1}(\gamma)&u(\gamma)\\ &\rho_{2}(\gamma)\\ \end{pmatrix},\ \gamma\in\Gamma,$
where $u:\Gamma\rightarrow\textup{Hom}(V_{2},V_{1})$ is a matrix valued function. The group $\rho_{1}(\Gamma)$ preserves the properly convex domain $\Omega_{0}\cap\mathbb{P}(V_{1})$ of $\mathbb{P}(V_{1})$. By [17, 39], there exists a closed $\rho_{1}(\Gamma)$-invariant properly convex domain $\Omega_{1}\subset\mathbb{P}(V_{1})$ and a $\rho_{1}(\Gamma)$-invariant closed convex subset $\mathcal{C}\subset\Omega_{1}$ such that $\rho_{1}(\Gamma)\backslash\mathcal{C}$ is compact. We fix a basepoint $x_{0}\in\mathcal{C}$ such that every point of $\mathcal{C}$ is within $d_{\Omega_{1}}$-distance $M>0$ from the orbit $\rho_{1}(\Gamma)\cdot x_{0}$. Let $g\in\Gamma$ and consider $x_{0},x_{1},\ldots,x_{k}\in[x_{0},gx_{0}]$ with $\frac{1}{2}\leqslant d_{\Omega}(x_{i},x_{i+1})\leqslant 1$. For every $0\leq i\leq k$, choose $g_{i}\in\Gamma$ such that $d_{\Omega}(\rho_{1}(g_{i})x_{0},x_{0})\leqslant M$, where $g_{0}=e$ and $g_{k}=g$. Now we define $\{h_{i}\}_{i=0}^{k+1}$ as follows: $h_{0}=e$, $h_{i}=g_{i-1}^{-1}g_{i}$, $1\leqslant i\leqslant k$ and $h_{k+1}=e$. Observe that $g=h_{1}\cdot\cdot\cdot h_{k}$ and a straightforward computation shows that
@@
$b\frac{\sigma_{1}(\rho_{1}(\gamma))^{1-\varepsilon}}{\sigma_{1}(\rho_{2}(\gamma))}\geq 1,\ \sigma_{1}(\rho_{1}(\gamma))\geqslant Ae^{ad_{\Omega_{1}}(\rho_{1}(\gamma)x_{0},x_{0})},\ \Big{|}\log\frac{\sigma_{1}(\rho_{1}(\gamma))}{\sigma_{d_{1}}(\rho_{1}(\gamma))}-2d_{\Omega_{1}}(\gamma x_{0},x_{0})\Big{|}\leqslant E.$ (4)
@@
$\big{|}\big{|}u(g)\big{|}\big{|}$ $\leqslant C_{R}\sum_{i=1}^{k-1}\sigma_{1}\big{(}\rho_{1}(h_{1}\cdot\cdot\cdot h_{i})\big{)}\cdot\sigma_{1}\big{(}\rho_{2}(h_{i+1}\cdot\cdot\cdot h_{k})\big{)}\leqslant bC_{R}\sum_{i=0}^{k-1}\frac{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))\sigma_{1}(\rho_{1}(w_{i}))}{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))^{\varepsilon}}$
$=bC_{R}\sum_{i=0}^{k-1}\frac{1}{\sigma_{1}(\rho_{1}(g^{-1}w_{i}))\sigma_{1}(\rho_{1}(w_{i}^{-1}))}\cdot\frac{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))}{\sigma_{d_{1}}(\rho_{1}(w_{i}^{-1}g))}\cdot\frac{\sigma_{1}(\rho_{1}(w_{i}))}{\sigma_{d_{1}}(\rho_{1}(w_{i}))}\cdot\frac{1}{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))^{\varepsilon}}$
$\leqslant bC_{R}\sum_{i=0}^{k-1}\frac{e^{2E}}{\sigma_{1}(\rho_{1}(g^{-1}))}\cdot e^{2d_{\Omega_{1}}(\rho_{1}(w_{i}^{-1}g)x_{0},x_{0})}\cdot e^{2d_{\Omega_{1}}(\rho_{1}(w_{i})x_{0},x_{0})}\cdot\big{(}A^{-\varepsilon}e^{-a\varepsilon|w_{i}^{-1}g|_{\Gamma}}\big{)}$
$=\frac{bC_{R}e^{2E}}{\sigma_{1}(\rho_{1}(g^{-1}))A^{\varepsilon}}\sum_{i=0}^{k-1}e^{2d_{\Omega}(\rho(w_{i})x_{0},\rho(g)x_{0})}\cdot e^{2d_{\Omega}(\rho(x_{i})x_{0},x_{0})}\cdot e^{-a\varepsilon d_{\Omega}(\rho_{1}(w_{i}^{-1}g)x_{0},x_{0})}$
@@
$\sigma_{1}(\rho^{ss}(g))\leqslant\sigma_{1}(\rho(g))\leqslant L\sigma_{1}(\rho^{ss}(g)),$
hence (2) holds true for some $C_{\rho}>0$, depending only on $\rho$. The highest weight of the $\theta$-compatible representation $\tau$ is of the form $\chi_{\tau}=\sum_{\alpha\in\theta}n_{a}\omega_{\alpha}$, $n_{\alpha}>0$, and since there is $D>0$ such that $\langle\omega_{\alpha},\mu(\rho(g))-\mu(\rho^{ss}(g))\rangle\geq-D$ for every $g\in\Gamma$. Thus, we have
$\langle\chi_{\tau},\mu(\rho(g))-\mu(\rho^{ss}(g))\rangle\geq n_{\alpha}\langle\omega_{\alpha},\mu(\rho(g))-\mu(\rho^{ss}(g))\rangle-C\sum_{\beta\in\theta\smallsetminus\{\alpha\}}n_{\beta},$
and in particular, for every $g\in\Gamma$ we conclude that
$\big{\langle}\omega_{\alpha},\mu(\rho(g))-\mu(\rho^{ss}(g))\big{\rangle}\leq\frac{1}{n_{\alpha}}\Bigg{(}C_{\rho}+C\sum_{\beta\in\theta\smallsetminus\{\alpha\}}n_{\beta}\Bigg{)}.$

### v4

@@ 2.7. Convex cocompact groups
([19, Prop. 10.1]) Let $\Omega$ be a properly convex domain of $\mathbb{P}(\mathbb{R}^{d})$. For any $x_{0}\in\Omega$, there exists $\kappa>0$ such that for any $g\in\textup{Aut}(\Omega)$,
$\frac{1}{2}\log\frac{\sigma_{1}(g)}{\sigma_{d}(g)}\geqslant d_{\Omega}(gx_{0},x_{0})-\kappa.$
Let $\Gamma$ be a subgroup of $\mathsf{PGL}_{d}(\mathbb{R})$ preserving a properly convex domain $\Omega$. By using the previous proposition we can control the Gromov product with respect to $\varepsilon_{1}\in\mathfrak{a}^{\ast}$ as follows.
@@
By Theorem 2.5 there exists a finite subset $F$ of $\Gamma$ and $M>0$ such that for every $\gamma\in\Gamma$ there exists $f\in F$ such that $\log\frac{\ell_{1}(\gamma f)}{\ell_{d}(\gamma f)}\geq\log\frac{\sigma_{1}(\gamma))}{\sigma_{d}(\gamma)}-M$. The translation length of an isometry $g\in\textup{Aut}(\Omega)$ is exactly $\frac{1}{2}\log\frac{\ell_{1}(g)}{\ell_{d}(g)}$, see [17, Prop. 2.1]. In particular, if $\gamma\in\Gamma$ and $f\in F$ are as previously, we have that
@@
Then, by Proposition 2.7 and (2.7), we obtain $L>0$ such that
[…]
@@
The following lemma follows immediately from [19, Thm. 1.4] and [44, Thm. 1.27] and is used to pass from a $P_{1}$-Anosov representation to a convex cocompact action in some projective space.
@@
Let $V_{d}$ be the vector space of $d\times d$-symmetric matrices and $\mathsf{S}_{d}:\mathsf{GL}_{d}(\mathbb{R})\rightarrow\mathsf{GL}(V_{d})$ be the representation defined as follows $\mathsf{S}_{d}(g)X=gXg^{t}$ for $g\in\mathsf{GL}_{d}(\mathbb{R})$ and $X\in V_{d}$. For every $P_{1}$-Anosov representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, the representation $\mathsf{S}_{d}\circ\rho$ is $P_{1}$-Anosov and $\mathsf{S}_{d}(\rho(\Gamma))$ is a strongly convex cocompact subgroup of $\mathsf{GL}(V_{d})$.
Given two representations $\rho_{1}:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ and $\rho_{2}:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$, we say that $\rho_{1}$ uniformly dominates $\rho_{2}$ if there is $0<\epsilon<1$ with the property that for every $\gamma\in\Gamma$,
$(1-\epsilon)\log\ell_{1}(\rho_{1}(\gamma))\geqslant\log\ell_{1}(\rho_{2}(\gamma)).$
We will also need the following lemma for the proof of Proposition 1.12, which allows us to control the Cartan projection of an Anosov representation $\rho$ in terms of the Cartan projection of a semisimplification $\rho^{ss}$ (of $\rho$). We expect that this fact follows by the techniques of Guichard–Wienhard in [25, §5] showing that Anosov representations have strong proximality properties.
@@
Let $\Gamma$ be a word hyperbolic group, $G$ a real semisimple Lie group and $\theta\subset\Delta$ a subset of simple restricted roots of $G$. Suppose $\psi:\Gamma\rightarrow G$ is a $P_{\theta}$-Anosov representation with semisimplification $\psi^{ss}:\Gamma\rightarrow G$. There is a constant $C_{\psi}>0$, depending only on $\psi$, such that for every $\gamma\in\Gamma$,
$\max_{\alpha\in\theta}\big|\omega_{\alpha}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\big|\leqslant C_{\psi}.$
@@
Let us first observe that for any linear representation $\phi:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ and any semisimplification $\phi^{ss}$ of $\phi$, by Theorem 2.5, there exists a constant $M>0$, depending only on $\phi$, such that
$\log\sigma_{1}(\phi(\gamma))\geq\log\sigma_{1}(\phi^{ss}(\gamma))-M$
for every $\gamma\in\Gamma$. Moreover, by [42, Thm. 7.2], for every $\alpha\in\theta$, there exists $N_{\alpha}>0$ such that $N_{\alpha}\omega_{\alpha}$ is the highest weight of an irreducible proximal representation $\tau_{\alpha}:G\rightarrow\mathsf{GL}_{m}(\mathbb{R})$. In particular, by the definition of the highest weight, there exists $M^{\prime}>0$, depending only on $\tau_{\alpha}$, such that
$\forall h\in G,\ \big|\log\sigma_{1}(\tau_{\alpha}(h))-N_{\alpha}\omega_{\alpha}(\mu(h))\big|\leq M^{\prime}.$
Given the representation $\psi:\Gamma\rightarrow G$, by the previous two facts, there is $C_{1}>0$, depending only on $\psi$ and $G$, such that for every $\gamma\in\Gamma$,
$\omega_{\alpha}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\geq-C_{1}.$ (2)
Now we prove that there is $D>0$ such that for every $\gamma\in\Gamma$,
$\omega_{\alpha}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\leqslant D.$
By Proposition 2.4, we may compose $\psi$ with an irreducible representation $\tau_{\theta}:G\rightarrow\mathsf{GL}_{n}(\mathbb{R})$ such that $\rho:=\tau_{\theta}\circ\psi$ and its semisimplification $\rho^{ss}=\tau_{\theta}\circ\psi^{ss}$ are $P_{1}$-Anosov. Clearly if $\rho$ is semisimple then the bound follows by Theorem 2.5. Hence, we continue by assuming that $\rho$ is not semisimple (hence non irreducible) and preserves some proper subspace of $\mathbb{R}^{n}$. Up to composing $\rho$ with the representation $\mathsf{S}_{n}$ from Lemma 2.10, we may further assume that $\rho(\Gamma)$ and the dual $\rho^{\ast}(\Gamma)$ preserve (possibly different) properly convex domains in $\mathbb{P}(\mathbb{R}^{n})$. Moreover, up to conjugating $\rho$, and possibly considering the dual representation of this conjugate, we may assume $\rho(\Gamma)$ preserves a properly convex domain $\Omega_{0}\subset\mathbb{P}(\mathbb{R}^{n})$ and there is a decomposition $\mathbb{R}^{n}=V_{1}\oplus\cdots\oplus V_{\ell}$ such that
$\rho=\begin{pmatrix}\rho_{1}&\ast&\ast\\ 0&\ddots&\ast\\ 0&0&\rho_{\ell}\end{pmatrix},\ \rho^{ss}=\begin{pmatrix}\rho_{1}&0&0\\ 0&\ddots&0\\ 0&0&\rho_{\ell}\end{pmatrix}$
where $\{\rho_{i}:\Gamma\rightarrow\mathsf{GL}(V_{i})\}_{i=1}^{\ell}$ are irreducible representations and $\rho_{1}$ is the restriction of $\rho^{ss}$ on the image of its limit map $\xi_{\rho^{ss}}^{+}:\partial_{\infty}\Gamma\rightarrow\mathbb{P}(\mathbb{R}^{n})$. By the definition of $V_{1}$, since the attracting fixed point of $\rho^{ss}(\gamma)$ lies in $V_{1}$ for every infinite order element $\gamma\in\Gamma$, the restriction of $\rho_{1}$ of $\rho^{ss}$ on $V_{1}$ uniformly dominates $\rho_{i}$ for every $2\leqslant i\leqslant\ell.$
@@
$\rho(\gamma)=\begin{pmatrix}\rho_{1}(\gamma)&u(\gamma)\\ 0&\rho_{2}(\gamma)\\ \end{pmatrix},\ \gamma\in\Gamma$
where $u:\Gamma\rightarrow\textup{Hom}(V_{2},V_{1})$ is an appropriate matrix valued function. The group $\rho_{1}(\Gamma)$ preserves the properly convex domain $\Omega_{0}\cap\mathbb{P}(V_{1})$ of $\mathbb{P}(V_{1})$. By [19, 44], there exists a closed $\rho_{1}(\Gamma)$-invariant properly convex domain $\Omega_{1}\subset\mathbb{P}(V_{1})$ and a $\rho_{1}(\Gamma)$-invariant closed convex subset $\mathcal{C}\subset\Omega_{1}$ such that $\rho_{1}(\Gamma)\backslash\mathcal{C}$ is compact. We fix a basepoint $x_{0}\in\mathcal{C}$ such that every point of $\mathcal{C}$ is within $d_{\Omega_{1}}$-distance $M>0$ from the orbit $\rho_{1}(\Gamma)\cdot x_{0}$. Let $g\in\Gamma$ and consider $x_{0},x_{1},\ldots,x_{k}\in[x_{0},gx_{0}]$ with $\frac{1}{2}\leqslant d_{\Omega}(x_{i},x_{i+1})\leqslant 1$. For every $0\leq i\leq k$, choose $g_{i}\in\Gamma$ such that $d_{\Omega}(\rho_{1}(g_{i})x_{0},x_{0})\leqslant M$, where $g_{0}=e$ and $g_{k}=g$. Now we define $\{h_{i}\}_{i=0}^{k+1}$ as follows: $h_{0}=e$, $h_{i}=g_{i-1}^{-1}g_{i}$, $1\leqslant i\leqslant k$ and $h_{k+1}=e$. Observe that $g=h_{1}\cdot\cdot\cdot h_{k}$ and a straightforward computation shows that
@@
$b\frac{\sigma_{1}(\rho_{1}(\gamma))^{1-\varepsilon}}{\sigma_{1}(\rho_{2}(\gamma))}\geq 1,\ \sigma_{1}(\rho_{1}(\gamma))\geqslant Ae^{ad_{\Omega_{1}}(\rho_{1}(\gamma)x_{0},x_{0})},\ \Bigg|\log\frac{\sigma_{1}(\rho_{1}(\gamma))}{\sigma_{d_{1}}(\rho_{1}(\gamma))}-2d_{\Omega_{1}}(\gamma x_{0},x_{0})\Bigg|\leqslant E.$ (4)
@@
[…]
$=bC_{R}\sum_{i=0}^{k-1}\Bigg(\frac{1}{\sigma_{1}(\rho_{1}(g^{-1}w_{i}))\sigma_{1}(\rho_{1}(w_{i}^{-1}))}\cdot\frac{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))}{\sigma_{d_{1}}(\rho_{1}(w_{i}^{-1}g))}\cdot\frac{\sigma_{1}(\rho_{1}(w_{i}))}{\sigma_{d_{1}}(\rho_{1}(w_{i}))}\cdot\frac{1}{\sigma_{1}(\rho_{1}(w_{i}^{-1}g))^{\varepsilon}}\Bigg)$
$\leqslant bC_{R}\sum_{i=0}^{k-1}\Bigg(\frac{e^{2E}}{\sigma_{1}(\rho_{1}(g^{-1}))}\cdot e^{2d_{\Omega_{1}}(\rho_{1}(w_{i}^{-1}g)x_{0},x_{0})}\cdot e^{2d_{\Omega_{1}}(\rho_{1}(w_{i})x_{0},x_{0})}\cdot A^{-\varepsilon}e^{-a\varepsilon|w_{i}^{-1}g|_{\Gamma}}\Bigg)$
$=\frac{bC_{R}e^{2E}}{\sigma_{1}(\rho_{1}(g^{-1}))A^{\varepsilon}}\sum_{i=0}^{k-1}\Bigg(e^{2d_{\Omega}(\rho(w_{i})x_{0},\rho(g)x_{0})}\cdot e^{2d_{\Omega}(\rho(x_{i})x_{0},x_{0})}\cdot e^{-a\varepsilon d_{\Omega}(\rho_{1}(w_{i}^{-1}g)x_{0},x_{0})}\Bigg)$
@@
$\sigma_{1}(\rho^{ss}(g))\leqslant\sigma_{1}(\rho(g))\leqslant L\sigma_{1}(\rho^{ss}(g)).$ (6)
Now recall that $\rho=\tau_{\theta}\circ\psi$ and $\rho^{ss}=\tau_{\theta}\circ\psi^{ss}$, where $\psi:\Gamma\rightarrow G$ is a $\theta$-Anosov representation, $\psi^{ss}$ is a semisimplification of $\psi$ and $\tau_{\theta}:G\rightarrow\mathsf{GL}_{n}(\mathbb{R})$ is a $\theta$-compatible representation. The highest weight of $\tau_{\theta}$ is of the form $\chi_{\tau_{\theta}}=\sum_{\alpha\in\theta}n_{a}\omega_{\alpha}$, where $n_{\alpha}>0$ and $\omega_{\alpha}$ is the fundamental weight with respect to $\alpha$ (e.g. see the discussion in [24, Subsec. 3.2]). By the definition of $\chi_{\tau_{\theta}}$, we may choose $D_{1}>0$, depending only on $\tau_{\theta}$, with the property for every $h\in G$,
$\Big|\log\sigma_{1}(\tau_{\theta}(h))-\chi_{\tau_{\theta}}(\mu(h))\Big|\leq D_{1}.$
Since $\rho^{ss}=\tau_{\theta}\circ\psi^{ss}$, by (6), there is $D>0$ such that for every $g\in\Gamma$,
$\Big|\chi_{\tau_{\theta}}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\Big|\leq D.$
By (2) there is $C_{1}>0$, depending only on $\psi$, with $\omega_{\alpha}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\geq-C_{1}$ for every $\gamma\in\Gamma$, thus, for every $\alpha\in\theta$
$\chi_{\tau_{\theta}}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)\geq n_{\alpha}\omega_{\alpha}\big(\mu(\psi(\gamma))-\mu(\psi^{ss}(\gamma))\big)-C_{1}\sum_{\beta\in\theta\smallsetminus\{\alpha\}}n_{\beta}.$
In particular, for every $\gamma\in\Gamma$ we conclude that
$\omega_{\alpha}\big(\mu(\rho(\gamma))-\mu(\rho^{ss}(\gamma))\big)\leq\frac{1}{n_{\alpha}}\Bigg(D+C\sum_{\beta\in\theta\smallsetminus\{\alpha\}}n_{\beta}\Bigg).$
This concludes the proof of the lemma. ∎